OpenStudy (anonymous):

How many solutions does the equation −4y + 4y + 2 = 2 have?

OpenStudy (anonymous):

many

alones (alones):

You have answer choices?

OpenStudy (anonymous):

How many solutions does the equation −4y + 4y + 2 = 2 have? Two None One Infinitely many

OpenStudy (anonymous):

yah answer choices

OpenStudy (anonymous):

many

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

AloneS can you help?

alones (alones):

Sure i can try ;)

OpenStudy (booklover40):

There is 1 answer

OpenStudy (photon336):

probably just one solution, well technically it's just 2 = 2

OpenStudy (anonymous):

yah

alones (alones):

after photons the last step subtract 2 from both sides. 2−2=2−2 0=0

OpenStudy (anonymous):